Assuming a constant pressure gradient, by what factor must a blood vessel decrease in radius if the blood flow is reduced by 65.4 percent?

THE ANSWER IS NOT .81
Answer #1

The ratio is, 요=A (젓 100 34.6 34.6 100 1: 0.346-ב = 0.77 1. The radius has to be reduced by about 23%The ratio is, 요=A (젓 100 34.6 34.6 100 1: 0.346-ב = 0.77 1. The radius has to be reduced by about 23%++